OpenStudy (unklerhaukus):

y' = (x-y+2)/(x-y+1) find y(x)

OpenStudy (unklerhaukus):

i think i made a mistake somewhere

hartnn (hartnn):

just a typo there, integral w+1 instead of integral w

OpenStudy (unklerhaukus):

2

OpenStudy (unklerhaukus):

thanks @hartnn , im still not getting what wolfram get

hartnn (hartnn):

wolf is getting something more complicated than your aolution

OpenStudy (unklerhaukus):

yeash its just the i factor

hartnn (hartnn):

isn't your solution acceptable ? i do not see any mistake there

hartnn (hartnn):

oh wait!

OpenStudy (unklerhaukus):

shouldn't there be two solution sets?

OpenStudy (unklerhaukus):

w^2/2+w =-x+c w^2+2w =2(-x+c) w^2+2w+1=2(-x+c)+1 (w+1)^2 =-2x+2c+1 (w+1)^2 =2(-x+C) w+1 =√(2(-x+C))

OpenStudy (unklerhaukus):

w^2/2+w =-x-c w^2+2w =-2(x+c) w^2+2w+1=-2(x+c)+1 (w+1)^2 =-2x-2c+1 (w+1)^2 =-2(x+C) w+1 =√(-2(x+C)) x-y+1 =i√2√(x+C) \[ y =x-i√2\sqrt{x+C}+1\]

hartnn (hartnn):

\((w+1)^2= 2(C-x) \\ w+1 = \pm \sqrt 2 \sqrt {C-x} \\ y =x+1 \mp \sqrt{2}\sqrt {C-x}\) if we purposely bring the imaginary term, \(y =x+1 \mp i\sqrt{2}\sqrt {x-C}\)

hartnn (hartnn):

if you look at the wolf solution you will notice that the \(\huge x^2\) term inside sqrt cancels out! giving you the form of \(\sqrt {x-C}\) only! you're correct :)

hartnn (hartnn):

\(\sqrt {c_1 -2 (-x^2/2) -2(-2x)-(x^2+2x+1)} \\ \sqrt {c_1 +x^2 +4x-x^2-2x-1} \\ \sqrt {c_2+2x} \\ \sqrt 2 \sqrt {x+c_2/2} = \sqrt 2 \sqrt {x-C} \)

OpenStudy (unklerhaukus):

hmm yeah, I wasn't really expecting \(i\) in the solution, none other problem in this book have solutions with \(i\), is there some other method or some other form, that might appear simpler?

OpenStudy (unklerhaukus):

or do i have to find a domain for C to make the solution real valued only

hartnn (hartnn):

i don't see how we get a more simplified solution than \( y =x+1 \mp \sqrt{2}\sqrt {C-x}\) also, i never really cared about the domain of constant...is it required to show that in your work ?

OpenStudy (unklerhaukus):

i dont know, i just wasn't expecting a complex answer,

OpenStudy (unklerhaukus):

ah!, if i solve it like an exact equation i get a different looking answer,

OpenStudy (unklerhaukus):

if i was going to use the method i tried first , maybe i should leave the solution implicitly , as \[(x-y)^2=-2x+c\]

OpenStudy (unklerhaukus):

Using the exact method , i got f(x,y) = x^2/2-xy+2x+y^2/2-y = k (x-y)^2 +4x-2y = k

OpenStudy (unklerhaukus):

if i was going to use the method i tried first , maybe i should leave the solution implicitly , as \[(x-y+1)^2=-2x+c\] ___ Using the exact method , i got \[ f(x,y) = x^2/2-xy+2x+y^2/2-y = k\]\[\qquad\qquad\qquad\qquad (x-y)^2 +4x-2y = k\] And these plot the same thing, so they are the same,

hartnn (hartnn):

yes, but the question specifically asked to fin y

OpenStudy (unklerhaukus):

sorry about that , thanks again!

hartnn (hartnn):

no problem! welcome ^_^